NaiNoah:

im cooked arent i The height h in feet of an object shot straight up with initial velocity v in feet per second is given by h = −16t2 + vt + c, where c is the initial height of the object above the ground. A model rocket is shot vertically up from a height of 6 feet above the ground with an initial velocity of 22 feet per second. Will it reach a height of 10 feet? Identify the correct explanation for your answer.

Yanaisswaggy:

i got a brain ingury reading this

Yanaisswaggy:

butttttt

Yanaisswaggy:

c = 4 ft v = 48 ft/s We want to know if the height will be 40 at a certain time t. h = 40 Plug in these values of c, v, and h into the equation. 40 = -16t2 + 48t + 4 Subtract 40 on both sides of equation. 0 = -16t2 + 48t - 36 We can factor out the 2 on the right side of the equation to simplify it. h = 2(-8t2 + 24t + 13) The discriminant is the square-root part of the quadratic formula. The value under the square-root is b2 - 4ac where: a = -8 b = 24 c = 13 242 - 4(-8)(13) = 576 + 416 = 992 The value under the square-root is positive. Therefore, when we solve for t, we will have a real value of time. In conclusion, the object will reach a height of 40 ft when we solve for that time.

NaiNoah:

@yanaisswaggy wrote:
c = 4 ft v = 48 ft/s We want to know if the height will be 40 at a certain time t. h = 40 Plug in these values of c, v, and h into the equation. 40 = -16t2 + 48t + 4 Subtract 40 on both sides of equation. 0 = -16t2 + 48t - 36 We can factor out the 2 on the right side of the equation to simplify it. h = 2(-8t2 + 24t + 13) The discriminant is the square-root part of the quadratic formula. The value under the square-root is b2 - 4ac where: a = -8 b = 24 c = 13 242 - 4(-8)(13) = 576 + 416 = 992 The value under the square-root is positive. Therefore, when we solve for t, we will have a real value of time. In conclusion, the object will reach a height of 40 ft when we solve for that time.
then tell me this No; The discriminant is negative, so the rocket will not reach a height of 10 feet. Yes; The discriminant is zero, so the rocket will reach a height of 10 feet. Yes; The discriminant is positive, so the rocket will reach a height of 10 feet. No; The discriminant is positive so the rocket will reach a height of 10 feet. these are the option choices

Yanaisswaggy:

pretty sure its no, but not 100% sure

NaiNoah:

@yanaisswaggy wrote:
pretty sure its no, but not 100% sure
my grade depends on you, ill let yk when I get the grade back

Yanaisswaggy:

OMG

Yanaisswaggy:

ur stupid

Yanaisswaggy:

im not smart

Yanaisswaggy:

but i tried my best

NaiNoah:

@yanaisswaggy wrote:
but i tried my best
if I fail my grade goes from a A to a B

Yanaisswaggy:

omg

Yanaisswaggy:

im so sorry if its bad

Yanaisswaggy:

im going to kill myself if its bad

NaiNoah:

@yanaisswaggy wrote:
im going to kill myself if its bad
don't say that lmao

Yanaisswaggy:

its finna be my fault

NaiNoah:

@yanaisswaggy wrote:
its finna be my fault
it'll be fine

Yanaisswaggy:

:\

velmalovesshaggy145:

projectile motion?

velmalovesshaggy145:

parametric equations?

NaiNoah:

@velmalovesshaggy145 wrote:
parametric equations?
The Discriminant

velmalovesshaggy145:

oh its a similar concept i believe.

surjithayer:

\[h=-16t^2+22t+6\] when h=10 ft \[-16t^2+22t+6=10\] \[-16t^2+22t-4=0\] divide by -2 \[8t^2-11t+2=0\] \[discriminant=(-11)^2-4\times 8 \times2=121-64=57>0\] so roots are real Hence it can reach 10 ft above the ground.